OpenStudy (anonymous):

what number must you add to the polynomial below to complete the square x^2-4x

jimthompson5910 (jim_thompson5910):

Take half of the x term to get (-4/2) = -2 Then square that result to get (-2)^2 = (-2)*(-2) = 4 So you must add 4 to complete the square

OpenStudy (anonymous):

thank-you!!!

jimthompson5910 (jim_thompson5910):

This will make x^2-4x into x^2-4x+4 which factors to (x-2)^2 and this is a perfect square.

jimthompson5910 (jim_thompson5910):

You're welcome